Refer to the following financial statements for Crosby Corporation CROSBY CORPORATION Income Statement For the Year Ended December 31, 20X2 3,900,000 2,470,000 1,430,000 618,000 290,000 522,000 es Cost of goods sold Selling and administrative expense Depreciation expense Operating income Interest expense 434,900 127,000 307,900 10,000 297,900 Eamings before taxes Taxes Eamings after taxes Preferred stock dividends Earnings available to common stockholders Shares outstanding Earnings per share Statement of Retained Earnings For the Year Ended December 31, 20X2 Retained earnings, balance, January 1, 20X2 $1,079,600 297,900 170,000 $1,207,500 Add: Earnings available to common stockholders, 20X2 Deduct: Cash dividends declared and paid in 20X2 Retained earnings, balance, December 31, 20X2 Comparative Balance Sheets For 20X1 and 20X2 Year-End Current assets $ 102,000 507,000 698,000 60,300 $ 1,367,300 97,000 $48,500 Accounts receivable (net) Inventory Prepaid expenses 754,000 37,900 $ 1,440,400 Total current assets Investments (long-term securities) Gross plant and equipment Less: Accumulated depreciation Net plant and equipment $2,630,000 1,370,000 $3,280,000 1,00040001620,000 1,260,000 $2,724,300 Total assets $ 3,148,500 Liabilities and Stockholders Equity Current liabilities: Accounts payable tes payable Accrued expenses Total current liabilities 553,000 75,700 949,700 553,000 53,000 $ 1,189,000 Long-term liabilities Bonds payable, 20X2 Total liabilities 105,000 $1,054,700 162,000 $1,351,000 Stockholders equity: Preferred stock, $100 par value Common stock, $1 par value Capital paid in excess of par Retained earnings Total stockholders equity 90,000 150,000 150,000 350,000 1,079,600 1,207,500 $ 1,797,500 $ 3,148,500 Total liabilities and stockholders equity $2,724,300a. Prepare a statement of cash flows for the Crosby Corporation: (Amounts to be deducted should be indicated with parentheses or a minus sign.) CROSBY CORPORATION Statement of Cash Flows For the Year Ended December 31, 20x2 Cash flows from operating activities: Adjustments to determine cashflow from operating activities: Total adjustments Net cash flows from operating activities Cash flows from investing activities Net cash flows from investing activities Cash flows from financing activities Net cash flows from financing activities b. Compute the book value per common share for both 20X1 and 20X2 for the Crosby Corporation. (Round your answers to 2 decimals places.) Book value 20X1 20X2c. If the market value of a share of common stock is 3.3 times book value for 20X1, what is the firms P/E ratio for 20X2 vs. 20X1? (Do not round intermediate calculations. Round your final answer to 2 decimal places.) P/E ratio times
Answer #1

Cash flow from operating Activities Net Income Adjustment to determine cashflow from oper act Add back depreciaiton Increase in account receivable Increase in inventory Decrease in prepaid expenses increase in account payable decrease in accrued expenses Total adjustments Net Cash flows from opearting activities Cash flow from Investing Activities Decrease in investments increase in plant and equipment Net Cash flow from investing activities Amount A B $307,900 Cash flow from operating Activitics Net Income Adjustment to determine cashflow from oper act 307900 S 290,000 $93,000 $56,000 $ 22,400 5 262,000 S-22,700 5 Add back depreciaiton Increase in account receivable Increase in inventory Decreas increase in account payable decrease in accruede 507000-600000 -698000-754000 -60300-37900 #583000-32 1000 -53000-75700 e in prepaid expenses S 402,700 11 Total adjustments $ 710,600 12 Net Cash flows from opearting activitles SUMID5:01 E3 E11 Cash flow from Investing Activities Decrease in investments increase in plant and equipment Net Cash flow from investing activities 97000-88100 -1260000-1620000 16 S 360,000 D15+D16 $ 351,100 1 Cash flow from Financine Activities Increase in bonds payab;e -162000-105000 Cash flow from Financing Activities eferred stock dividend paid 10000 170000 S 57,000 S 10,000 S-170,000 common stock dividend paid preferred stock dividend paid common stock dividend paid Net Cash Provided by financing activities 23Net Cash Provided by financing activities SUMID20:D22 123,000 25 Nat Increase in cash flows E12+E17+E23

b)20x1=(total equity-preferred stock)/common shares outstanding
=(1669600-10000)/150000
=11.06
20x2=(1797500-10000)/150000
=11.92

c) market value=3.3 times book value 20x1=3.3*11.06=36.50
PE ration in 20x2=market price/EPS in 20x2
=36.50/1.99
=18.34

> How do you get the common stock dividents paid? i cant find it.

Salvador456 Sun, Jan 30, 2022 9:38 PM